OpenStudy (anonymous):

MEDAL FASTEST ANSWER! can someone teach me how to do this please:) What is the value of the expression 2/4 to the -3 power

OpenStudy (jdoe0001):

hmmm woops \(\bf \left( \cfrac{2}{4}\right)^{-3}?\) rather

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

doesn't that equal 8?? right?

OpenStudy (anonymous):

my answer choices r

OpenStudy (anonymous):

128 256 1/6 1/32

OpenStudy (anonymous):

2 ---- -3 4

OpenStudy (anonymous):

thats the problem

OpenStudy (misty1212):

HI!!

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):



OpenStudy (misty1212):

first off \[\frac{2}{4}=\frac{1}{2}\] right?

OpenStudy (jdoe0001):

pfffft

OpenStudy (anonymous):

it equals 1/32

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

oh \[\huge \frac{2}{4^{-3}}\]?

OpenStudy (anonymous):

yeah

OpenStudy (misty1212):

ok negative exponent means flip it get \[\huge 2\times 4^3\]

OpenStudy (anonymous):

ok